# ABCD-LINK: Annotation Bootstrapping for Cross-Document Fine-Grained Links

### v1

@@ Abstract
Understanding fine-grained relations between documents is crucial for many application domains. However, the study of automated assistance is limited by the lack of efficient methods to create training and evaluation datasets of cross-document links. To address this, we introduce a new domain-agnostic framework for selecting a best-performing approach and annotating cross-document links in a new domain from scratch. We first generate and validate semi-synthetic datasets of interconnected documents. This data is used to perform automatic evaluation, producing a shortlist of best-performing linking approaches. These approaches are then used in an extensive human evaluation study, yielding performance estimates on natural text pairs. We apply our framework in two distinct domains – peer review and news – and show that combining retrieval models with LLMs achieves 78% link approval from human raters, more than doubling the precision of strong retrievers alone. Our framework enables systematic study of cross-document understanding across application scenarios, and the resulting novel datasets lay foundation for numerous cross-document tasks like media framing and peer review. We make the code, data, and annotation protocols openly available. [^1] https://github.com/UKPLab/arxiv2025-ABCD-Link
@@ 1 Introduction
Documents rarely exist in isolation. In many practical scenarios, understanding one document requires reasoning over its relationships with others. Fact-checkers might trace a claim to the specific sentences across multiple articles that support or contradict it (Thorne et al., 2018; Wadden et al., 2020; Chen et al., 2024b). Authors and meta-reviewers reading peer reviews need to connect reviewer comments to the paper to decide on the course of action (Kuznetsov et al., 2022; D’Arcy et al., 2024). News analysts might use paraphrased or ideologically reframed sentences across different outlets to gain insights into reporting bias and source alignment (Giorgi et al., 2023). Yet finding such fine-grained relations can be cognitively demanding, motivating the need for machine assistance. Following Kuznetsov et al. (2022), we call fine-grained relations between documents links, and the general task of identifying them linking. While automatic linking holds great potential, the progress has been limited due to the manual effort required to annotate and evaluate links, and the diversity of link types across domains. Thus, labeled corpora of links are scarce, preventing the development and evaluation of automated linking approaches at scale.
To move beyond the bottleneck of full manual annotation, we propose a general framework for bootstrapping sentence-level cross-document linking without the need for pre-existing labeled data (Figure 1). While links might be hard for humans to detect, linked documents are easy to generate with state of the art LLMs. Building upon this insight, we (1) propose a method to generate and validate semi-synthetic corpora of linked documents. We use this data to (2) automatically evaluate a wide range of approaches to arrive at a shortlist of best-performing linking approaches. Finally, we use the best-performing approaches to (3) assist humans in labeling natural, non-synthetic document pairs in a large-scale human-in-the-loop evaluation study. Our framework allows practitioners to efficiently select the best linking approach given a domain and link type, while producing manually labeled many-to-many cross-document linking datasets for downstream applications.
@@ 2 Related Work
Numerous NLP tasks study or rely on cross-document relations, including cross-document coreference (Miyabe et al., 2008; Cybulska and Vossen, 2014; Ravenscroft et al., 2021), fact-checking (Thorne et al., 2018; Wadden et al., 2020; Chen et al., 2024b), scholarly peer review analysis (Kuznetsov et al., 2022; D’Arcy et al., 2024), and citation/source detection (Syed et al., 2023; Liang et al., 2024). While prior work focuses on limited relation types and relies on manual labeling, we contribute a general, structured framework for bootstrapping linking from scratch in new domains and for new link types. We evaluate our framework in two domains: peer reviews and news, which target challenging and underrepresented forms of linking, such as subjective evaluation and ideological framing, as opposed to encyclopedic and factual relations found in Wikipedia (Feith et al., 2024).
@@
Human judgment is essential for evaluating NLP systems, yet it is difficult to elicit reliably. There is growing interest in structured human evaluation protocols, particularly for tasks involving model-assisted annotation or subjective decision-making. Prior work examined inter-annotator calibration over time (Uma et al., 2021), preference-based assessment of model outputs (Clark et al., 2021), and the influence of interface design on annotation quality (Klie et al., 2018). In the context of LLM evaluation, human-in-the-loop setups have proven essential for capturing fine-grained distinctions and improving label quality (Zhou et al., 2023). Here, we contribute a novel and efficient annotation protocol that combines candidate pre-selection with a ranking-based comparison for a reliable assessment of top-performing approaches.
@@ 3 Setup
To trade off between task complexity and utility, we focus on sentence-level linking, cast as a sentence pair classification task. Given two documents, the source document $A=\{a_{1},\ldots,a_{N}\}$ and the target document $B=\{b_{1},\ldots,b_{M}\}$, the goal is to predict for each cross-document sentence pair $(a_{i},b_{j})$ whether a relation $r$ holds between them. The definition accommodates many-to-many links, both directed (e.g. a review sentence criticizing a claim in a paper) and undirected (e.g. sentences in two new articles addressing the same fact). Linking can be seen as a special case of information retrieval, where $a_{i}$ serves as a query to retrieve relevant sentences from $B$. However, linking is constrained by a task-specific relationship holding between sentences, and might require documents $A$ and $B$ for contextualizing the linking decisions.
@@
Stat
NEWS
ECB+
NEWS
SYNTH
REVIEWS
SYNTH
REVIEWS
F1000
@@ 4.1 Generation
To match the number of links in REVIEWS-SYNTH, for NEWS-SYNTH we sample 346 news articles from the WikinewsSum dataset (Calizzano et al., 2022). To reduce noise in the documents, we clean each article using GPT-4o-mini to remove scraping artifacts. We then segment the text into sentences, and prompt the LLM to produce a related article on the same topic, written with a different tone or editorial perspective, and instructing the model to link sentences in the generated article to sentences in the original article, resulting in cross-document links. This simulates a use case in media analysis where the same events reported by different outlets are compared to detect bias and framing. See Appendix M for more details, prompt templates and example links for both datasets.
To contextualize our synthetic datasets and experimental results, we derive two additional datasets from existing cross-document corpora. NEWS-ECB is derived from the ECB+ corpus (Cybulska and Vossen, 2014) designed for cross-document coreference resolution. We repurpose this dataset for sentence-level linking by aligning sentence pairs across documents that share event mentions in the same cluster. REVIEWS-F1000 leverages the F1000RD dataset introduced by Kuznetsov et al. (2022), which links peer review sentences to their corresponding paper sentences. Although its annotation design restricts the choices of possible links, it is the most structurally similar resource to our task. Dataset conversion details are provided in Appendix A. Table 1 provides key statistics on the synthetic and converted datasets.
@@ 4.2 Validation
We ensure the quality of the data through automatic and manual inspection. We first compare synthetic and natural documents using document-level statistics: lexical diversity, subjectivity, and lexical complexity (Flesch-Kincaid Reading Ease score (Flesch, 1948)). In NEWS-SYNTH, synthetic articles are more lexically diverse, similarly subjective, and slightly more lexically complex than their natural counterparts. In contrast, REVIEWS-SYNTH synthetic reviews show lower lexical diversity and higher complexity. We attribute this to LLMs’ tendency to use formal or technical language within the prompt’s length constraints, whereas natural reviews are longer and can use simpler phrasing for communication. Subjectivity remains comparable across domains. Overall, synthetic documents align well with natural ones in most stylistic dimensions. See Appendix B for further details.
To further assess the quality of the synthetic data, we conduct a human validation study. For NEWS-SYNTH, we randomly selected 30 natural news articles and their corresponding synthetic counterparts. Four annotators rated each article on a five-point Likert scale for fluency, coherence, realism (i.e., could the article be written by a human), and specificity (i.e., does the article contain specific facts). Synthetic articles match fluency/coherence but score lower on realism and specificity, likely due to LLMs producing more general statements. While these limitations may affect journalistic authenticity, they do not hinder our linking task where coherence is more important, because it provides better context to reason over.
For REVIEWS-SYNTH, we randomly selected 30 synthetic reviews and retrieved natural reviews for the same papers from the NLPeer dataset. Three NLP PhD students rated each review on fluency, coherence, helpfulness (i.e., would this review be helpful for the paper’s authors), and specificity (i.e., does the review address concrete parts of the paper). Synthetic reviews were rated higher in fluency and coherence, and lower in helpfulness and specificity due to a lack of concrete suggestions and feedback. While not perfect, the results align with our goal of generating reviews that are coherent and linkable. See Appendix C for further details on the setup.
In summary, our results suggest that LLM-generated linked document pairs, while sometimes less specific or realistic, offer a practical, scalable way to create sentence-level links that can be used for automatic evaluation, to which we now turn.
Figure 2: R+LLM Setup.
@@
Table 2: Performance of retrieval models across synthetic and converted datasets. Metrics are: average F1 score across all cutoffs ($k\in\{1,3,5,7,10,20\}$) and recall at a fixed cutoff (R@10 for news, R@20 for reviews). The overall average F1 is computed across all datasets. Recall cutoffs are selected based on domain-specific document lengths to maximize retrieval coverage. Detailed per-dataset and per-cutoff results are provided in Appendix D.
@@ 5.1 Approaches
With our synthetic data, we can test a wide range of retrieval models in a zero-shot setup with the goal of shortlisting top-performing models for the annotation and human evaluation study. We test sparse models (BM25 (Robertson and Zaragoza, 2009), SPLADEv3 (Lassance et al., 2024), BGE-M3 (Chen et al., 2024a)), bi-encoders (all-mpnet-base-v2222https://huggingface.co/sentence-transformers/all-mpnet-base-v2, SFR-Embedding-Mistral (Meng et al., 2024), BGE-M3, Contriever (Izacard et al., 2022), Dragon+ (Lin et al., 2023)), and cross-encoders (ms-marco-MiniLM-L6-v2333https://huggingface.co/cross-encoder/ms-marco-MiniLM-L6-v2, BGE-M3). Each retriever ranks candidate sentences in the target document based on cosine similarity to a sentence from the source document. We convert ranks into binary decisions via a threshold: all sentences that appear within the top-$k$ are treated as links.
Next, we explore the use of LLMs to refine the retrieval results. We use the best-performing retrieval model to retrieve the top-$k$ candidate target sentences for each source sentence to reduce the search space. We choose the $k$ based on the performance of the retrieval models, and the length of the documents, to maximize the Recall@$k$. For NEWS domain datasets, we use top-10 results, and for PEER REVIEW domain, we use top-20. We then use the LLM to classify the source-target pairs as linked or not linked (Figure 2). We use LLMs for their zero-shot ability to adapt to different link types via prompting only, making them well-suited for domains where link semantics vary and labeled data for fine-tuning is unavailable. We further refer to this approach as R+LLM for brevity.
We investigate two prompting setups. In Pairwise, the LLM is given one source-target sentence pair at a time from the top-$k$ list to classify them independently. In Listwise, the LLM is given the source sentence and all $k$ target sentences at once, and classifies them in a single pass. In each setup, we prompt the LLM with both whole documents in four configurations: with a link description, with in-context examples, with both, and with no guidance. For evaluation, we use two open-source models, Phi-4 (14B) Abdin et al. (2024) and Qwen2.5 (32B) Yang et al. (2024), selected for their SOTA performance within their respective size classes. We also include GPT-4o [^4] Version gpt-4o-2024-08-06, as the SOTA closed-source non-reasoning model at the time of writing, to serve as a high-performance reference point. Prompt templates are given in Appendix F.
Figure 3: F1 scores across datasets using different prompt configurations and models (Phi-4, Qwen2.5, GPT4o). Results show a consistent trend across domains with listwise prompting outperforming pairwise prompting. Furthermore, combination prompts (tested listwise only) achieve the highest F1 scores on synthetic datasets, while performance gaps narrow on the converted datasets. These results showcase the interaction between prompt design, model capabilities, and dataset complexity.
@@ 5.2 Results
Table 2 summarizes the performance of all retrieval models across synthetic and converted datasets. The sparse models perform slightly worse overall. Dense models such as SFR, Dragon+, and BGE-M3 (Dense) demonstrate stronger F1 scores across all datasets. Among them, Dragon+ achieves the highest average F1 across datasets and cutoffs, outperforming all other models. This ranking is calculated by averaging the F1 scores computed at each cutoff ($k\in\{1,3,5,7,10,20\}$) for each dataset, followed by averaging across all datasets. While the performance gap between the top bi-encoders and cross-encoders is not large, due to higher inference costs of cross-encoders, along with Dragon+ consistently ranking among the strongest models, we select it as the base retriever for subsequent experiments.
@@
Building on the best retriever (Dragon+), we next experiment with using LLMs as classifiers to refine the results. Figure 3 presents the F1 scores of the different R+LLM setups. On the synthetic datasets (NEWS-SYNTH and REVIEWS-SYNTH), applying the LLM as a filter consistently improves performance, particularly for REVIEWS-SYNTH, which benefits from the LLM’s ability to capture nuanced feedback and critique relations. On the converted datasets (NEWS-ECB+ and REVIEWS-F1000), however, the gains are marginal. We attribute this to the inherent mismatch between these datasets and our linking task: NEWS-ECB+ focuses on coreference links, while REVIEWS-F1000 is derived from F1000RD whose papers are on average longer than papers in REVIEWS-SYNTH, and more importantly due to some of its domains (biology, medicine) key evidence appears in figures, limiting LLM effectiveness in our text-only setup. Nonetheless, LLMs still provide added value in capturing nuanced relations (e.g. support, critique, or clarification) that are implicit or span multiple lines of reasoning. Thus, while both domains benefit from R+LLM, the nature of their linking tasks determines how large the performance gap over the baseline is. An error analysis of common failure patterns of R+LLM is detailed in Appendix E.
@@
To evaluate the necessity of retrieval, we ablate the retriever and use LLMs alone for classification. We use the aforementioned best prompt setup and instruct the models to classify every sentence in the target document as linked or not, details are provided in Appendix G. Across all datasets, the LLM-only approach underperforms the R+LLM approach, confirming that retrieval is crucial for narrowing the search space and filtering distractors. The performance gap is especially large on the longer more complex peer review documents. For example, GPT-4o reaches only 33.42 and 39.22 F1 on REVIEWS-SYNTH and REVIEWS-F1000, compared to 41.42 and 59.18 with R+LLM. Even on the shorter news articles, the LLM-only setup falls short (e.g., 74.27 vs. 84.02 on NEWS-SYNTH). While LLMs can capture some cross-document signals without retrieval, their high inference cost makes them inefficient as standalone solutions compared to their performance, making retrieval as a first step essential for both efficiency and accuracy.

### v2

@@ Abstract
Understanding fine-grained links between documents is crucial for many applications, yet progress is limited by the lack of efficient methods for data curation. To address this limitation, we introduce a domain-agnostic framework for bootstrapping sentence-level cross-document links from scratch. Our approach (1) generates and validates semi-synthetic datasets of linked documents, (2) uses these datasets to benchmark and shortlist the best-performing linking approaches, and (3) applies the shortlisted methods in large-scale human-in-the-loop annotation of natural text pairs. We apply the framework in two distinct domains – peer review and news – and show that combining retrieval models with LLMs achieves a 73% human approval rate for suggested links, more than doubling the acceptance of strong retrievers alone. Our framework allows users to produce novel datasets that enable systematic study of cross-document understanding, supporting downstream tasks such as media framing analysis and peer review assessment. All code, data, and annotation protocols are released to facilitate future research. [^1] https://github.com/UKPLab/eacl2026-abcd-link
@@ 1 Introduction
Documents rarely exist in isolation. In many practical scenarios, understanding one document requires reasoning over its relationships with others. Fact-checkers might trace a claim to the specific sentences across multiple articles that support or contradict it (Thorne et al., 2018; Wadden et al., 2020; Chen et al., 2024b). Authors and meta-reviewers reading peer reviews need to connect reviewer comments to the paper to decide on the course of action (Kuznetsov et al., 2022; D’Arcy et al., 2024). News analysts might use paraphrased or ideologically reframed sentences across different outlets to gain insights into reporting bias and source alignment (Giorgi et al., 2023). Yet finding such fine-grained relations can be cognitively demanding, motivating the need for machine assistance. While tasks like cross-document coreference resolution offer some support, cross-document relations are much more diverse, necessitating the development of general-purpose approaches to cross-document analysis.
Following Kuznetsov et al. (2022), we call fine-grained relations between documents links, and the general task of identifying them linking. While automatic linking holds great potential, the progress has been limited due to the manual effort required to annotate and evaluate links, and the diversity of link types across domains. Thus, labeled corpora of links are scarce, preventing the development and evaluation of automated linking approaches at scale.
To move beyond the bottleneck of full manual annotation, we propose a general framework for bootstrapping sentence-level cross-document linking without the need for pre-existing labeled data (Figure 1). While links might be hard for humans to detect, linked documents are easier to generate with state of the art LLMs. Building upon this insight, we (1) propose a method to generate and validate semi-synthetic corpora of linked documents. We use this data to (2) automatically evaluate a wide range of approaches to arrive at a shortlist of best-performing linking approaches. Finally, we use the best-performing approaches to (3) assist humans in labeling natural, non-synthetic document pairs in a large-scale human-in-the-loop evaluation study. Our framework allows practitioners to efficiently select the best linking approach given a domain and link type, while producing manually labeled many-to-many cross-document linking datasets for downstream applications.
@@ 2 Related Work
Numerous NLP tasks study or rely on cross-document relations, including cross-document coreference resolution (CDCR) (Miyabe et al., 2008; Cybulska and Vossen, 2014; Ravenscroft et al., 2021), fact-checking (Thorne et al., 2018; Wadden et al., 2020; Chen et al., 2024b), scholarly peer review analysis (Kuznetsov et al., 2022; D’Arcy et al., 2024), citation/source detection (Syed et al., 2023; Liang et al., 2024), and cross-document question answering Lin et al. (2025). While prior work focuses on limited relation types and relies on manual labeling, we contribute a general, structured framework for bootstrapping linking from scratch in new domains and for new link types. We evaluate our framework in two domains: peer reviews and news, which target challenging and underrepresented forms of linking, such as subjective evaluation and ideological framing, as opposed to encyclopedic and factual relations found in Wikipedia (Feith et al., 2024). Crucially, unlike CDCR where the main goal is to resolve and cluster entity and event mentions, linking focuses on detecting a broad spectrum of sentence-level relationships, such as paraphrases, quotes and implicit references.
@@
Human judgment is essential for evaluating NLP systems, yet difficult to elicit reliably. There is growing interest in structured human evaluation protocols, particularly for tasks involving model-assisted annotation or subjective decision-making. Prior work examined inter-annotator calibration over time (Uma et al., 2021), preference-based assessment of model outputs (Clark et al., 2021), and the influence of interface design on annotation quality (Klie et al., 2018). In the context of LLM evaluation, human-in-the-loop setups have proven essential for capturing fine-grained distinctions and improving label quality (Zhou et al., 2023; Min et al., 2025). Here, we contribute a novel and efficient annotation protocol that combines candidate pre-selection with a ranking-based comparison for a reliable assessment of top-performing approaches.
@@ 3 Setup
We focus on sentence-level linking cast as a sentence pair classification task, representing a trade-off between task complexity and utility: a more narrow unit of analysis restricts the available context, while the links between broader units can be reconstructed by aggregating consecutive sentences. [^2] While broader argumentative or narrative relations that span multiple consecutive sentences do exist, in practice, many-to-many relations naturally emerge as clusters of consecutive sentence-level links. We provide an exploratory analysis of such clusters in Appendix A Given two documents, the source document $A=\{a_{1},\ldots,a_{N}\}$ and the target document $B=\{b_{1},\ldots,b_{M}\}$, the goal is to predict for each cross-document sentence pair $(a_{i},b_{j})$ whether a relation $r$ holds between them. The definition accommodates many-to-many links, both directed (e.g. a review sentence criticizing a claim in a paper) and undirected (e.g. sentences in two new articles addressing the same fact). Linking can be seen as a special case of information retrieval, where $a_{i}$ serves as a query to retrieve relevant sentences from $B$. However, linking is constrained by a task-specific relationship holding between sentences, and might require documents $A$ and $B$ for contextualizing the linking decisions.
@@
The framework is designed to be domain-agnostic in that it does not depend on any domain-specific characteristics like writing style, relationship type or the presence of explicit references such as citations. To demonstrate this, we apply the framework in two distinct practical scenarios: in REVIEWS, we link peer reviews to their papers, and in NEWS, we link pairs of news articles. While the framework itself aims to be domain-agnostic, the particular link types (and cutoff values, see Section 5) are flexible and defined per domain. In REVIEWS, we define a link as a sentence in a review that comments on, critiques, or supports a specific sentence in the corresponding paper. In NEWS, a link connects two sentences from different articles that convey the same or closely related factual content, typically reflect semantic equivalence or paraphrastic overlap, even if the framing or tone differs between the two sources.
@@
4 Synthetic data
4.1 Generation
Stat NEWS ECB+ NEWS SYNTH REVIEWS SYNTH REVIEWS F1000
@@ 4.1 Generation
To match the number of links in REVIEWS-SYNTH, for NEWS-SYNTH we sample 346 news articles from the WikinewsSum dataset (Calizzano et al., 2022). To reduce noise in the documents, we clean each article using GPT-4o-mini to remove scraping artifacts. We then segment the text into sentences, and prompt the LLM to produce a related article on the same topic, written with a different tone or editorial perspective, and instructing the model to link sentences in the generated article to sentences in the original article, resulting in cross-document links. This simulates a use case in media analysis where the same events reported by different outlets are compared to detect bias and framing. See Appendix P for more details, prompt templates and example links for both datasets.
To contextualize our synthetic datasets and experimental results, we derive two additional datasets from existing cross-document corpora. NEWS-ECB+ is derived from the ECB+ corpus (Cybulska and Vossen, 2014) designed for cross-document coreference resolution. We repurpose this dataset for sentence-level linking by aligning sentence pairs across documents that share event mentions in the same cluster. REVIEWS-F1000 leverages the F1000RD dataset introduced by Kuznetsov et al. (2022), which links peer review sentences to their corresponding paper sentences. Although its annotation design restricts the choices of possible links, it is the most structurally similar resource to our task. Dataset conversion details are provided in Appendix B. Table 1 provides key statistics on the synthetic and converted datasets.
Figure 2: R+LLM Setup.
@@ 4.2 Validation
We ensure the quality of the synthetic data, and explore potential biases and hallucinations in machine-generated data, through automatic and manual inspection. At the document level, we compare lexical diversity, subjectivity, and complexity (Flesch-Kincaid score (Flesch, 1948)). In NEWS-SYNTH, synthetic articles are more lexically diverse, similarly subjective, and slightly more lexically complex than naturally occurring articles. In REVIEWS-SYNTH, synthetic reviews are less diverse but more complex, reflecting LLMs’ preference for concise, formal phrasing. Subjectivity remains comparable across domains. Overall, synthetic documents align well with natural ones in most stylistic dimensions (more details in Appendix C)
@@
Table 2: Performance of retrieval models across synthetic and converted datasets. Metrics are: average F1 score across all cutoffs ($k\in\{1,3,5,7,10,20\}$) and recall at a fixed cutoff (R@10 for news, R@20 for reviews). The overall average F1 is computed across all datasets. Recall cutoffs are selected based on domain-specific document lengths to maximize retrieval coverage. Detailed per-dataset and per-cutoff results are provided in Appendix E.
To further assess the quality of the synthetic data, we conduct a human validation study. For NEWS-SYNTH, we randomly selected 30 natural news articles and their corresponding synthetic counterparts. Four annotators rated each article on a five-point Likert scale for fluency, coherence, realism (i.e., could the article be written by a human), and specificity (i.e., does the article contain specific facts). Synthetic articles match fluency/coherence but score lower on realism (as shown in Figure 14(a) in the Appendix) which we attribute to the models limitation in generating realistic sounding news articles. The synthetic data also shows lower specificity, likely due to LLMs producing more general statements compared to natural articles. While these limitations may affect journalistic authenticity, they do not hinder our linking task where coherence is more important, because it provides better context to reason over.
For REVIEWS-SYNTH, we randomly selected 30 synthetic reviews and retrieved natural reviews for the same papers from the NLPeer dataset. Three NLP PhD students rated each review on fluency, coherence, helpfulness (i.e., would this review be helpful for the paper’s authors), and specificity (i.e., does the review address concrete parts of the paper). Synthetic reviews were rated higher in fluency and coherence, and lower in helpfulness and specificity due to a lack of concrete suggestions and feedback. While not perfect, the results align with our goal of generating reviews that are coherent and linkable. See Appendix D for further details on the setup.
In summary, while LLM-generated linked document pairs are sometimes less specific or natural, they offer a practical, scalable way to create sentence-level links that can be used for automatic evaluation.
@@ 5.1 Approaches
Using our synthetic data, we benchmark a broad range of retrievers in a zero-shot setup to shortlist strong candidates for human evaluation. We test sparse models (BM25 (Robertson and Zaragoza, 2009), SPLADEv3 (Lassance et al., 2024), BGE-M3 (Chen et al., 2024a)), bi-encoders (all-mpnet-base-v2333sentence-transformers/all-mpnet-base-v2, SFR-Embedding-Mistral (Meng et al., 2024), BGE-M3, Contriever (Izacard et al., 2022), Dragon+ (Lin et al., 2023)), and cross-encoders (ms-marco-MiniLM-L6-v2444cross-encoder/ms-marco-MiniLM-L6-v2, BGE-M3). Each retriever ranks candidate sentences in the target document based on cosine similarity to a sentence from the source document. We convert ranks into binary decisions via a threshold: all sentences that appear within the top-$k$ are treated as links.
Figure 3: F1 scores across datasets using different prompt configurations and models (Phi-4, Qwen2.5, GPT-4o). Results show a consistent trend across domains with listwise prompting outperforming pairwise prompting. Furthermore, combination prompts (tested listwise only) achieve the highest F1 scores on synthetic datasets, while performance gaps narrow on the converted datasets. These results showcase the interaction between prompt design, model capabilities, and dataset complexity.
We then extend this setup with LLMs to refine candidate links. The best-performing retriever supplies the top-$k$ target sentences. We choose the $k$ based on the performance of the retrieval models, and the length of the documents, to maximize the ($k=10$ for NEWS domain, $k=20$ for REVIEWS domain). Next, the LLM classifies the source-target pairs as linked or not linked (Figure 2). We use LLMs for their zero-shot ability to adapt to different link types via prompting only, making them well-suited for domains where link semantics vary and labeled data for fine-tuning is unavailable. We further refer to this approach as R+LLM for brevity.
We investigate two prompting setups. Pairwise, where each source–target pair is judged independently, and Listwise, where all $k$ candidates are considered jointly in a single pass. In each setup, we prompt the LLM with both whole documents in four configurations: with a link description, with in-context examples, with both, and with no guidance. For evaluation, we use two open-source models, Phi-4 (14B) Abdin et al. (2024) and Qwen2.5 (32B) Yang et al. (2024), selected for their SOTA performance within their respective size classes. We also include GPT-4o [^5] Version gpt-4o-2024-08-06, as the SOTA closed-source non-reasoning model at the time of writing, to serve as a high-performance reference point. Prompt templates are given in Appendix I.
@@ 5.2 Results
Table 2 reports performance across synthetic and converted datasets. Sparse models perform weakest overall, while dense models (SFR, Dragon+, BGE-M3) demonstrate stronger F1 scores across all datasets. Overall, Dragon+ achieves the highest average F1 across datasets and cutoffs. This ranking is calculated by averaging the F1 scores computed at each cutoff ($k\in\{1,3,5,7,10,20\}$) for each dataset, followed by averaging across all datasets. While the performance gap to the cross-encoders model is not large, due to their higher inference costs, we select Dragon+ as the base retriever for subsequent experiments.
@@
Building on Dragon+, we next experiment with using LLMs as classifiers to refine the results. Figure 3 shows the F1 scores of the different R+LLM setups. On the synthetic datasets (NEWS-SYNTH and REVIEWS-SYNTH), applying the LLM as a filter leads to consistent improvements, particularly for REVIEWS-SYNTH, which benefits from the LLM’s ability to capture nuanced feedback and critique relations. On the converted datasets (NEWS-ECB+ and REVIEWS-F1000), however, the gains are marginal. We attribute this to the mismatch between these datasets and our linking task: NEWS-ECB+ focuses on coreference links, while REVIEWS-F1000 is derived from F1000RD whose papers are on average longer than papers in REVIEWS-SYNTH, and include domains (biology, medicine) where key evidence appears in figures, limiting LLM effectiveness in our text-only setup. Significance tests (Appendix F) support these observations. Overall, LLMs add value across domains by capturing relations that are implicit or span multiple lines of reasoning. Thus, while both domains benefit from R+LLM, the nature of their linking tasks determines how large the performance gap over the baseline is. Manual error analysis reveals that most failures stem from scattered or ambiguous source sentences, which make fine-grained linking difficult (more details in Appendix G). Finally, to verify that the improvements of R+LLM are not specific to Dragon+, we also tested the pipeline with other retrievers and observed similar performance gains (see Appendix H for details).
@@
To evaluate the necessity of retrieval, we ablate the retriever and use LLMs alone for classification. We use the aforementioned best prompt setup and instruct the models to classify every sentence in the target document as linked or not, (more details in Appendix J). As shown in Figure 4, across all datasets, the LLM-only approach underperforms the R+LLM approach, confirming that retrieval is crucial for narrowing the search space and filtering distractors. The performance gap is especially large on the longer more complex peer review documents. For example, GPT-4o reaches only 33.42 and 39.22 F1 on REVIEWS-SYNTH and REVIEWS-F1000, compared to 41.42 and 59.18 with R+LLM. Even on the shorter news articles, the LLM-only setup falls short (e.g., 74.27 vs. 84.02 on NEWS-SYNTH). While LLMs can capture some cross-document signals without retrieval, their high inference cost makes them inefficient as standalone solutions compared to their performance, making retrieval as a first step essential for both efficiency and accuracy.